solve 3/4(5+2x)+3is greater then or equal to 15
Answer by mananth(16946)

3/4(5+2x)+3>=15
add -3
3/4(5+2x)+3-3>=15-3
3/4(5+2x)>=12
multiply by 4/3
3/4*4/3(5+2x)>=4/3*12
(5+2x)>=16
add -5
+5-5+2x>=16-5
2x>=11
divide by 2
2x/2>=11/2
x>=11/2